OpenStudy (comrad):

A force of 18 pounds acts on an object while a force of 20 pounds acts on the same object at an angle of 120 with the first force. find the magnitude of the resultant force.

OpenStudy (comrad):

@dan815 Yo, I'm short on time, can you help me out?

OpenStudy (anonymous):

I think it would be 19.1 pounds

OpenStudy (michele_laino):

the situation of the problem can be represented as follows:  so the magnitude of the resultant force, is given by the theorem of Carnot: \[\Large R = \sqrt {{{18}^2} + {{20}^2} - \left( {2 \times 18 \times 20 \times \cos 60} \right)} = ...?\]